### v1

# A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation

Abstract
[…]
1 Introduction
High Dynamic Range (HDR) [4, 57] images capture a broader range of intensity values compared to Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels. From human vision perspective, LDR images are often not visually pleasing, while from computer vision applications perspective, they hold limited information.
Most commercial devices produce and display LDR content and some specialized hardware can support HDR, e.g., HDR cameras and sensors [37], which can capture images with more than $256$ intensity levels and HDR displays [12], which can display HDR intensity levels. Given that specialized hardware that captures HDR is expensive, researchers have been investigating approaches for accurate HDR reconstruction from LDR (i.e., inverse tone-mapping [78]) using conventional non-learning methods and modern data-driven approaches. For display purposes, however, one often needs to tone-map [27] the HDR content to LDR to fit the intensity range supported by standard hardware.
Early attempts on HDR reconstruction address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [62] or low lighting conditions [84, 47] and compression [11]. Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [83], media and entertainment [29], gaming, mixed reality and novel view synthesis [72, 53, 75, 31], as well as medical imaging [31].
Early data-driven methods utilize Convolutional Neural Networks (CNN) [74], Transformers [73, 52] and Generative Adversarial Networks (GAN) [67, 63]. More recent approaches are based on Diffusion Models [15, 23], Neural Radiance Fields (NeRF) [31], and Gaussian Splatting [75]. Some methods use multi-exposed LDR [6, 71, 18, 48, 51, 45, 24, 90, 25, 69], while others use single-exposed LDR [5, 9, 88] paired with an HDR for supervised training.
Although these methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 51, 19] for training. Consequently, the quality of the state-of-the-art data-driven HDR reconstruction methods depends on the quality of the available paired {LDR,HDR} datasets. There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [49], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction process [23, 22, 52, 77].
To address this gap, we propose a method that leverages cycle consistency [89, 14] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing CLIP [66] embeddings and loss based on semantic segmentation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
[…]
  • We propose novel generators based on modified U-Net architecture [70] that incorporates ConvLSTM-based feedback mechanism [30, 50] to mitigate visual artifacts in the HDR reconstruction (cf., Section 3.1).
  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR pairs, while maximizing the difference between LDR and other {LDR,HDR} pairs (cf., Section 3.1).
  • We propose a novel loss function based on the Mean Intersection of Union (mIOU) metric to further ensure semantic consistency between the LDR and reconstructed HDR (cf., Section 3.2).
  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (cf., Section 5).
2 Related Work
2.1 Non-Learning Approaches
The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [35]. This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image. Another approach is using histogram equalization [16] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image. Gradient domain manipulation [44] techniques enhance the granular details of an LDR image and expand its dynamic range. Pyramid-based image fusion [38] is a technique where image features are extracted into Laplacian pyramids [38] and then fused for the HDR. Some methods use Retinex-based [40, 58] tuning to effectively produce and display HDR images on consumer screens [40, 58]. Intensity Mapping Function based approaches [85, 17, 65] use transformations on the image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels. The methods in this category, although time-efficient, face some issues including ghosting effects, artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
2.2 Learning-Based Approaches
The methods in this category either use single-exposed LDR as input [71, 18, 48, 51, 45, 24, 90, 25] or alternatively, multi-exposed LDR [63, 9, 69, 88, 6]. Barua et al. [5] harnessed the concept of histogram equalization of input LDR images, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR image. Cao et al. [10] proposed a channel-decoupled kernel-based approach which combines the output HDR with the output of another architecture in a pixel-wise fashion. Luzardo et al. [54] proposed a method to enhance the artistic intent of the reconstructed HDR by enhancing the peak brightness in the reconstruction process. In addition, recently, Neural Radiance Fields (NeRF) [60, 31, 59, 53], Diffusion Models [15, 23], and Gaussian Splatting [75] have also been used to improve the reconstruction, both in terms of space and time complexity.
Some methods use weakly-supervised [45], self-supervised [88] and unsupervised [62] approaches for HDR reconstruction. Le et al. [45] proposed an indirect approach for HDR reconstruction from multi-exposed LDR using weak supervision. The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [28] to obtain an HDR image. Zhang et al. [88] proposed a self-supervised approach for HDR reconstruction. The method requires multi-exposed LDR to learn the reconstruction network during training without the need for HDR counterparts. Nguyen et al. [62] proposed an unsupervised approach to recover image information from overexposed areas of LDR images. This approach does not require any HDR for supervision and instead uses pseudo ground truth images. Lee et al. [46] proposed a method with limited supervision for multi-exposed LDR generation consisting of a pair of networks that generate LDR with higher and lower exposure levels. Then the two exposure levels are combined to generate the HDR image.
GAN-Based Approaches. Generative Adversarial Networks (GAN) [21, 3] architectures are well-explored in this domain. Niu et al. [63] proposed a GAN-based method that can handle images with foreground motions by fusing multi-exposed LDR and extracting important information from the over/underexposed areas of the LDR. Raipurkar et al. [67] proposed a conditional GAN architecture that adds details to the saturated regions of the input LDR using a pre-trained segmentation model to extract exposure masks. Guo et al. [24] proposed a two-stage pipeline that extracts the over/underexposed features with high accuracy. GAN with attention mechanism first generates the missing information in those extreme exposure areas, and in the second stage, a CNN with multiple branches fuses the multi-exposed LDR from the previous stage to reconstruct the HDR. Nam et al. [61] proposed a GAN method that uses exposure values for conditional generation of multi-exposure stack that adapts well to varying color and brightness levels.
Li et al. [49] proposed a GAN-based method for unpaired multi-exposure LDR to HDR translation. The method introduces modified GAN loss and novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR. Wu et al. [82] proposed CycleGAN-like architecture for low-light image enhancement task using unpaired data for training. The method fuses the Retinex theory with CycleGAN concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR image.
Table 1: Summary of recent state-of-the-art methods. I/O: LDR used as input (SE: Single-exposed and ME: Multi-exposed), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), Unpaired: Uses unpaired data, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts, TM: Also performs tone-mapping i.e. HDR $\rightarrow$ LDR.
Method I/O O/P Unpaired Context Semantics Artifacts TM
PSENet [62] SE D
SingleHDR(W) [45] SE I
UPHDR-GAN [49] ME D
SelfHDR [88] ME I
KUNet [77] SE D
Ghost-free HDR [52] ME D
DITMO [23] SE I
Ours SE D
Semantic and Knowledge-Based Approaches. Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction. Wang et al. [77] proposed a method which approximates the inverse of the camera pipeline. Their knowledge-inspired block uses the image formation knowledge to address three tasks during HDR reconstruction: missing details recovery, image parameters adjustment, and image noise reduction. Goswami et al. [23] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera. The proposed method works in two stages - first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions. Liu et al. [52] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR. The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
Limitations. Our analysis reveals that methods based on single-exposed LDR fail to preserve image characteristics such as color hue and saturation, color contrast, and intensity brightness levels in the HDR. On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts, missing details and ghosting effects. While most methods produce excellent results in terms of visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images. Table 1 summarizes a comparison between recent state-of-the-art and our method in terms of various parameters.
Figure 1: Overview of the proposed method architecture (cf., Section 3.1) where $x$ and $y$ represent LDR and HDR images, respectively. The method is trained with five objectives: adversarial, cycle consistency, identity, contrastive, and semantic segmentation (cf., Section 3.2).
3 Method
[…]
3.1 Architecture Modules
We adopt a CycleGAN [89] architecture as the basis of our method depicted in Figure 1. The network includes two generators and two discriminators (cf., Supplementary for brief background on cycle consistency). Let us denote the two domains as $X$ for LDR images and $Y$ for HDR images. Furthermore, let us denote the generator used in the forward cycle that maps images from LDR to HDR as $G_{Y}$ and the one used in the backward cycle that maps images from HDR to LDR as $G_{X}$. Finally, let us denote the discriminator that discriminates between the reconstructed LDR and real LDR images as $D_{X}$ and the one that discriminates between the reconstructed HDR and real HDR images as $D_{Y}$. The images from the two domains can be represented as $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$. The data distribution of the two domains can be represented as $x\sim p_{\text{data}}(x)$ and $y\sim p_{\text{data}}(y)$.
Generators. The generators $G_{Y}$ and $G_{X}$ are based on U-Net architecture [70] that includes an encoder and a decoder block with skip connections from each level of the encoder to the decoder. In our generators, we propose a feedback mechanism [30, 86] between the encoder and decoder block. The rationale behind the feedback is to refine the features extracted from the encoder (during the first iteration of the feedback) to guide the decoder for better output image reconstruction over the rest of the iterations. Hence, the feedback block not only iterates over its own output but also re-runs the decoder situated ahead of it in each iteration while keeping the encoder frozen until the feedback iteration completes. The feedback is implemented with a ConvLSTM [30, 50] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (cf., Supplementary for visual depiction of the proposed generators and implementation details).
Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [49, 32] (cf., Supplementary for implementation details).
Encoders. We also introduce a CLIP embeddings encoder $E$. Specifically we use the pre-trained Vision Transformer (ViT) [2] CLIP encoder from [66] which helps in extracting image embeddings with both local and global semantic context. For the forward cycle consistency we use $E(x)$ and $E(y)$. Then we add the embeddings from the encoder and feed them back to the bottleneck layer of the $G_{Y}$ decoder. Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(x)$ and feed them back to the bottleneck layer of the $G_{X}$ decoder.
3.2 Loss Functions
We use five loss functions to train the proposed method. Two are standard CycleGAN loss functions, i.e., adversarial loss [20] and cycle consistency loss [89]. Inspired from [89] we also use an identity loss [76]. Finally, we introduce two novel loss functions for HDR reconstruction, i.e., contrastive and semantic segmentation loss.
The loss is calculated on tone-mapped versions of the reconstructed and real HDR images. The tone-mapping is performed on the basis of the $\mu$-law [34]. This is done to avoid the high intensity pixels of HDR images that can distort the loss calculation. The tone-mapping operator $T$ can be represented as follows:
$T(y_{j})=\frac{log(1+\mu y_{j})}{log(1+\mu)},$ (1)
where the amount of compression $\mu$ is 5000 following [39].
Adversarial Loss. We apply adversarial loss to both mappings. The mapping from LDR to HDR domain, i.e., $G_{Y}:X\rightarrow Y$ with the discriminator $D_{Y}$, can be expressed as:
$\begin{split}\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\log D_{Y}(y)\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log(1-D_{Y}(G_{Y}(x)))\right],\end{split}$ (2)
where $G_{Y}$ generates images that look similar to images from domain $Y$ while $D_{Y}$ distinguishes between generated samples $\hat{y}$ and real samples $y$. $G_{Y}$ aims to minimize this objective against $D_{Y}$ that aims to maximize it, i.e., $\min_{G_{Y}}\max_{D_{Y}}\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)$.
The mapping from HDR to LDR domain, i.e., $G_{X}:Y\rightarrow X$ with the discriminator $D_{X}$, can be expressed as:
$\begin{split}\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)=\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log D_{X}(x)\right]+\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\log(1-D_{X}(G_{X}(y)))\right],\end{split}$ (3)
where $G_{X}$ generates images that look similar to images from domain $X$ while $D_{X}$ distinguishes between generated samples $\hat{x}$ and real samples $x$. Similar to above, $G_{X}$ aims to minimize this objective against $D_{X}$ that aims to maximize it, i.e., $\min_{G_{X}}\max_{D_{X}}\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)$.
Cycle Consistency Loss. The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each another. Hence, we also incorporate a cycle consistency loss to prevent mutual contradiction of the learned mappings $G_{Y}$ and $G_{X}$. For each image $x_{i}$ from the LDR domain, the cycle of reconstruction i.e., from LDR to HDR and then back to LDR must result back in the original image $x_{i}$. Hence, we can define the forward cycle consistency as: $x_{i}\rightarrow G_{Y}(x_{i})\rightarrow G_{X}(G_{Y}(x_{i}))\approx x_{i}$. Similarly, the backward cycle consistency can be represented as: $y_{j}\rightarrow G_{X}(y_{j})\rightarrow G_{Y}(G_{X}(y_{j}))\approx y_{j}$. The loss can be formulated as:
$\begin{split}\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})=\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(G_{Y}(x))-x\|_{1}\right]+\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(G_{X}(y))-y\|_{1}\right],\end{split}$ (4)
where $\|\cdot\|_{1}$ is the L1 norm (cf., Supplementary for visual depiction of the cycle consistency loss).
Identity Loss. For LDR $\leftrightarrow$ HDR translation tasks, we also find that adversarial and cycle consistency formulations alone cannot preserve the color and hue information. This is due to incorrect mapping of color shades from LDR to HDR domains by the generators stemming from the underlying difference in dynamic ranges. Therefore, we force the generators to replicate an identity mapping by providing target domain images. This loss can be expressed as:
$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (5)
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [8]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(y)$ we first define the cosine similarity between them as follows:
$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (6)
[…]
$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (7)
where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs. This loss is calculated in both the forward and backward cycles with the input and output being swapped. This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (cf., Supplementary for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection of Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [8]. We choose equalized images instead of original LDR because segmentation on low-light or extreme brightness images does not yield good results. We use Segment Anything (SAM) [42] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images (cf., Supplementary for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as:
$\text{IoU}_{c}=\frac{x_{c}\cap y_{c}}{x_{c}\cup y_{c}},$ (8)
where $x_{c}$ and $y_{c}$ represent the segmentations for class $c$ in image $x$ and $y$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$. The mean IoU over all segmentation classes can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (9)
We represent the semantic segmentation loss as:
$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (10)
$\mathcal{L}_{\text{full}}=\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}$ (11)
Final Loss. The full objective is expressed in Eq. (11), where $\lambda$ scales the relative importance of the cycle consistency and identity loss. $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [89]. $\alpha$ and $\beta$ are weights for the contrastive and semantic segmentation loss, respectively. Both values are set to $2$ (cf., Supplementary for ablation experiments for $\alpha$ and $\beta$).
4 Experiments
4.1 Implementation
The proposed method is implemented in PyTorch [^1] https://pytorch.org/ on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD. We used Adam optimizer [41] to train the models for $170$ epochs. Batch size of $1$ is expected to work well in cycle consistency models [89] however, since we are using contrastive loss on positive and negative pairs, the batch size is set to $4$. We used learning rate of $4\times 10^{-4}$ for the generators and $2\times 10^{-4}$ for the discriminators. The learning rate is kept constant for the first $100$ epochs and subsequently set to decay linearly to 00. For all models, we resized the input images to $512\times 512$. The HDR images used in the text are tone-mapped with Reinhard’s operator [68].
4.2 Datasets
For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [13], NTIRE [64], and HDR-Synth & HDR-Real [51] paired datasets. HDRTV has $1235$ samples for training and $117$ for testing. The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples. HDR-Synth & HDR-Real dataset consists of $20537$ samples. Other paired datasets used in our evaluations are DrTMO [19] ($1043$ samples), Kalantari [36] ($89$ samples), HDR-Eye [43] (46 samples), and LDR-HDR Pair [33] ($176$ samples). We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity. We perform $80/20$ split for training and testing unless specified otherwise. For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR. For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [8] for datasets with only single-exposed LDR images.
4.3 Metrics
We use four metrics to report the results. High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [55] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model. Structural Similarity Index Measure (SSIM) [79, 80, 81] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information. Peak Signal-to-Noise Ratio (PSNR) [26] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^2] PSNR is computed on $\mu$-law based tone-mapped HDR images.. Learned Perceptual Image Patch Similarity (LPIPS) [87] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge based contributions. This metric aligns with the perceptual judgement of humans.
Table 2: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with supervised (gray) and unsupervised/weakly-supervised/self-supervised (black) learning methods trained on the paired datasets HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51].
[…]
HDRCNN [18] 20.12 0.724 0.422 60.12
DrTMO [19] 20.43 0.711 0.412 60.81
ExpandNet [56] 21.44 0.854 0.451 58.12
FHDR [39] 20.84 0.841 0.307 62.45
SingleHDR [51] 21.01 0.851 0.402 64.21
Two-stage [1] 34.29 0.856 0.287 61.78
HDR-GAN [63] 30.11 0.912 0.411 65.01
KUNet [77] 37.21 0.974 0.051 62.14
Ghost-free [52] 40.32 0.966 0.043 63.02
ArtHDR-Net [6] 37.12 0.961 0.321 63.43
HistoHDR-Net [5] 38.21 0.968 0.301 65.15
UPHDR-GAN [49] 39.98 0.965 0.044 63.54
PSENet [62] 27.35 0.856 0.274 62.89
SingleHDR(W) [45] 30.79 0.952 0.055 62.95
SelfHDR [88] 39.51 0.972 0.037 64.77
Ours 40.01 0.979 0.021 68.41
Table 3: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with UPHDR-GAN trained on unpaired datasets.
[…]
UPHDR-GAN [49] 42.51 0.980 0.043 66.13
Ours 43.92 0.989 0.020 69.32
5 Results
5.1 HDR Reconstruction
For an extensive comparison we choose a combination of methods that utilize different learning paradigms including, HDRCNN [18], DrTMO [19], ExpandNet [56], FHDR [39], SingleHDR [51], Two-stage [1], KUNet [77], HistoHDR-Net [5], HDR-GAN [63], ArtHDR-Net [6], Ghost-free HDR [52], SelfHDR [88], UPHDR-GAN [49], SingleHDR(W) [45], and PSENet [62]. These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised/weakly-supervised/self-supervised, unpaired learning, and knowledge-conditioned.
Table 4: LDR reconstruction (tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with the state-of-the-art one-mapping operators.
[…]
$\mu$-law operator [34] 28.68 0.891 0.311
Reinhard’s operator [68] 30.12 0.903 0.342
Photomatix [28] 32.11 0.922 0.136
Ours 31.71 0.941 0.092
Table 2 summarizes the results of evaluation on HDRTV [13], NTIRE [64], HDR-Synth & HDR-Real [51] paired datasets, the most widely used datasets by the state-of-the-art (cf., Supplementary for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $4$ unsupervised/weakly-supervised/self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics (cf., Supplementary for visual depiction of the quantitative HDR reconstruction results). It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion using only unpaired samples (cf., Section 3).
To further demonstrate the strength of our approach we extended the train set used in the previous experiment with: 1) The HDR images from DrTMO [19], Kalantari [36], HDR-Eye [43], LDR-HDR Pair [33], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images. We kept the test set identical to the previous experiment in order to compare with the state-of-the-art paired data based methods. Given that the only other method that supports this type of unpaired train data is UPHDR-GAN, we perform a direct comparison with that approach in Table 3. The results demonstrate that our method outperforms UPHDR-GAN and sets new state-of-the-art on all metrics (cf., Supplementary for more unpaired data experiments).
Table 5: Quantitative evaluation of our feedback U-Net-based generators for HDR reconstruction. For evaluation we use the dataset employed by SingleHDR(W), i.e., DrTMO [19]. The models marked with “(mod)” use feedback U-Net-based generators and the rest use the same U-Net generator as SingleHDR(W).
[…]
SingleHDR(W) [45] 30.79 0.952 0.055 62.95
SingleHDR(W) (mod) 30.03 0.961 0.036 63.91
Ours 35.11 0.969 0.029 67.88
Ours (mod) 43.01 0.988 0.023 68.21
Table 6: Architecture ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets. Uo: The U-Net generator used in SingleHDR(W) [45], Uf: The proposed U-Net generator, E: CLIP embedding encoder.
[…]
CycleGAN [89] 21.34 0.849 0.351 61.55
Uo 25.13 0.851 0.341 61.79
Uf 24.07 0.879 0.126 63.16
Uf+E 29.19 0.885 0.122 64.21
Uf+E+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
Uf+E+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
Uo+E+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 35.15 0.953 0.042 67.18
Ours 40.01 0.979 0.021 68.41
5.2 LDR Reconstruction
For a complete evaluation, we also report results for LDR image reconstruction in Table 4 and compare with the state-of-the-art tone-mapping operators $\mu$-law [34], Reinhard’s [68], and Photomatix [28]. The test data is identical to the one used in the HDR reconstruction experiments. The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS. In terms of PSNR, it is second best after Photomatix [68] (cf., Supplementary for more LDR reconstruction experiments).
5.3 Ablation Results
Generators. We propose a novel generator that modifies the U-Net architecture by introducing a feedback mechanism (cf., Section 3.1). The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult. To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [45] method with our U-Net generator. We see that leads to improvement in all metrics as shown in Table 5. The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W) (cf., Supplementary for qualitative results in support of this quantitative evaluation).
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the CLIP embedding encoder and feedback U-Net generators (cf., Section 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Table 6 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W) [45]. The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2. The fourth row lists the results when we add the CLIP embedding encoder to the architecture. This results in an improvement in all metrics. The fifth and sixth rows depict the improvements after adding contrastive and semantic segmentation objectives in the architecture (cf., Section 3.2). The seventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). Finally, the last row is the full architecture with all components which achieves the best performance on all metrics.
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Table 7 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{id}}$). In the third and forth rows we add the contrastive and semantic segmentation loss. This results in an improvement in all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$). Finally, adding all loss functions achieves the best results on all the metrics.
Table 7: Loss functions ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
$\mathcal{L}_{\text{GAN}}$+$\mathcal{L}_{\text{cyc}}$ 24.55 0.878 0.125 62.11
+$\mathcal{L}_{\text{id}}$ 29.19 0.885 0.122 64.21
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 40.01 0.979 0.021 68.41
We also perform separate tests on contrastive and semantic segmentation loss to examine the contribution of histogram-equalized LDR compared to original LDR images. We summarize these results in Tables 8 and 9. In both cases the loss calculated between the histogram-equalized versions of the LDR (instead of original the LDR) and the recnstructed HDR is better. This is due to the fact that histogram-equalized LDR can reveal semantic information in extremely over/underexposed areas of the original LDR images which in turn improves the extracted CLIP embeddings for $\mathcal{L}_{\text{con}}$ and segmentation maps for $\mathcal{L}_{\text{sem}}$.
Table 8: Contrastive loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
LDR 39.12 0.966 0.030 67.11
LDRHist 40.01 0.979 0.021 68.41
Table 9: Segmentation loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
LDR 38.89 0.956 0.033 66.92
LDRHist 40.01 0.979 0.021 68.41
6 Conclusions
Most current state-of-the-art methods for HDR reconstruction (inverse tone-mapping) require high-quality paired {LDR,HDR} datasets for model training. Consequently, the quality of the state-of-the-art in inverse tone-mapping depends on the quality of the available paired {LDR,HDR} datasets. Additionally, only a few approaches utilize semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images.
[…]

### v2

@@
-A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation
+LLM-HDR: Bridging LLM-based Perception and Self-Supervision for Unpaired LDR-to-HDR Image Reconstruction
 Abstract
 […]
+Figure 1: Qualitative comparison of the proposed LLM-HDR method and state-of-the-art method SelfHDR [116] (ICLR’24). We see, our method handles the overexposed portions in the sky more realistically.
 1 Introduction
-High Dynamic Range (HDR) [4, 57] images capture a broader range of intensity values compared to Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels.
-From human vision perspective, LDR images are often not visually pleasing, while from computer vision applications perspective, they hold limited information.
-Most commercial devices produce and display LDR content and some specialized hardware can support HDR, e.g., HDR cameras and sensors [37], which can capture images with more than $256$ intensity levels and HDR displays [12], which can display HDR intensity levels.
+High Dynamic Range (HDR) [3, 70] images capture a wider range of intensity values compared to their Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels.
+From human vision perspective, LDR images are often not visually pleasing, while from the computer/robot vision applications perspective, they hold limited information.
+Most commercial devices produce and display LDR content, and some specialized hardware can support HDR, e.g., HDR cameras and sensors [45], which can capture images with more than $256$ intensity levels, and HDR displays [14], which can display HDR intensity levels.
 Given that specialized hardware that captures HDR is expensive, researchers have been investigating approaches for accurate HDR reconstruction from LDR (i.e., inverse tone-mapping [98]) using conventional non-learning methods and modern data-driven approaches.
 For display purposes, however, one often needs to tone-map [33] the HDR content to LDR to fit the intensity range supported by standard hardware.
-Early attempts on HDR reconstruction address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [62] or low lighting conditions [84, 47] and compression [11].
-Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [83], media and entertainment [29], gaming, mixed reality and novel view synthesis [72, 53, 75, 31], as well as medical imaging [31].
+Early HDR reconstruction methods address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [76] or low lighting conditions [109, 57] and compression [13].
+Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [103], media and entertainment [35], gaming, mixed reality, and novel view synthesis [88, 66, 91, 37], as well as medical imaging [37].
 Early data-driven methods utilize Convolutional Neural Networks (CNN) [90], Transformers [89, 65] and Generative Adversarial Networks (GAN) [82, 77].
-More recent approaches are based on Diffusion Models [15, 23], Neural Radiance Fields (NeRF) [31], and Gaussian Splatting [75].
-Some methods use multi-exposed LDR [6, 71, 18, 48, 51, 45, 24, 90, 25, 69], while others use single-exposed LDR [5, 9, 88] paired with an HDR for supervised training.
-Although these methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 51, 19] for training.
+Recent approaches for HDR image and/or 3D HDR scene reconstruction are based on Diffusion Models [19, 28, 8, 95], Neural Radiance Fields (NeRF) [37], and Gaussian Splatting [91].
+Some use multi-exposed [5, 87, 22, 60, 64, 55, 29, 118, 30, 85], while others use single-exposed LDR [6, 11, 116] paired with HDR images for supervised training.
+Large Language Models (LLM) [75] and Vision-Language Models (VLM) [114, 105] have been widely used in many vision tasks, e.g., image generation [92], deepfake detection [41], semantic segmentation [97], image relighting [15], and adverse weather condition removal [108].
+However, the human-defined knowledge of such models has not been utilized for LDR-to-HDR image reconstruction.
+Although current methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 64, 23] for training.
 Consequently, the quality of the state-of-the-art data-driven HDR reconstruction methods depends on the quality of the available paired {LDR,HDR} datasets.
-There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [49], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction process [23, 22, 52, 77].
-To address this gap, we propose a method that leverages cycle consistency [89, 14] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR.
-In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing CLIP [66] embeddings and loss based on semantic segmentation.
+There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [61, 94], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction [28, 27, 65, 96].
+To address this gap, we propose LLM-HDR, a method that leverages LLM-driven cycle consistency [117, 17] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR.
+In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing Contrastive Language-Image Pretraining (CLIP) [81] embeddings and loss based on semantic segmentation.
+Moreover, LLM-based artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation.
 The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images.
@@
 […]
-  • We propose novel generators based on modified U-Net architecture [70] that incorporates ConvLSTM-based feedback mechanism [30, 50] to mitigate visual artifacts in the HDR reconstruction (cf., Section 3.1).
-  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR pairs, while maximizing the difference between LDR and other {LDR,HDR} pairs (cf., Section 3.1).
-  • We propose a novel loss function based on the Mean Intersection of Union (mIOU) metric to further ensure semantic consistency between the LDR and reconstructed HDR (cf., Section 3.2).
-  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (cf., Section 5).
+  • We leverage LLM-based loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs. 3.1 and 3.2).
+  • We propose a novel generator based on a modified U-Net architecture [86] that incorporates ConvLSTM-based artifact-aware feedback mechanism [36, 63] and exposure-aware skip connections to mitigate visual artifacts in the HDR reconstruction (Sec.
+3.1).
+  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR image pairs, while maximizing the difference between LDR and other {LDR,HDR} image pairs (Sec.
+3.1).
+  • We propose a novel loss function based on the Mean Intersection over Union (mIoU) metric to further ensure semantic consistency between the LDR and reconstructed HDR images (Sec. 3.2).
+  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (Sec. 5).
 2 Related Work
-2.1 Non-Learning Approaches
-The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [35].
-This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image.
-Another approach is using histogram equalization [16] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image.
-Gradient domain manipulation [44] techniques enhance the granular details of an LDR image and expand its dynamic range.
-Pyramid-based image fusion [38] is a technique where image features are extracted into Laplacian pyramids [38] and then fused for the HDR.
-Some methods use Retinex-based [40, 58] tuning to effectively produce and display HDR images on consumer screens [40, 58].
-Intensity Mapping Function based approaches [85, 17, 65] use transformations on the image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels.
-The methods in this category, although time-efficient, face some issues including ghosting effects, artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
-2.2 Learning-Based Approaches
-The methods in this category either use single-exposed LDR as input [71, 18, 48, 51, 45, 24, 90, 25] or alternatively, multi-exposed LDR [63, 9, 69, 88, 6].
+Learning-based approaches use either single-exposed LDR as input [87, 22, 60, 64, 55, 29, 118, 30] or alternatively, multi-exposed LDR [77, 11, 85, 116, 5].
+Some of the multi-exposed methods use novel feature fusion or aggregation methods [111, 104, 107].
 Barua et al.
-[5] harnessed the concept of histogram equalization of input LDR images, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR image.
+[6] harnessed the concept of histogram equalization of input LDR, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR.
 Cao et al.
@@
 [67] proposed a method to enhance the artistic intent of the reconstructed HDR by enhancing the peak brightness in the reconstruction process.
-In addition, recently, Neural Radiance Fields (NeRF) [60, 31, 59, 53], Diffusion Models [15, 23], and Gaussian Splatting [75] have also been used to improve the reconstruction, both in terms of space and time complexity.
-Some methods use weakly-supervised [45], self-supervised [88] and unsupervised [62] approaches for HDR reconstruction.
+In addition, recently, Neural Radiance Fields [72, 37, 73, 66], Diffusion Models [19, 28, 95], and Gaussian Splatting [91] have also been used to improve the reconstruction of HDR content.
+Some methods use weakly-supervised [55], self-supervised [116] and unsupervised [76, 94] approaches for HDR reconstruction.
 Le et al.
 [55] proposed an indirect approach for HDR reconstruction from multi-exposed LDR using weak supervision.
-The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [28] to obtain an HDR image.
+The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [34] to obtain an HDR.
 Zhang et al.
@@
 Nguyen et al.
-[62] proposed an unsupervised approach to recover image information from overexposed areas of LDR images.
-This approach does not require any HDR for supervision and instead uses pseudo ground truth images.
+[76] proposed an unsupervised approach to recover image information from overexposed areas of LDR.
+This approach does not require any HDR for supervision and instead uses pseudo-ground truth images.
 Lee et al.
 [56] proposed a method with limited supervision for multi-exposed LDR generation consisting of a pair of networks that generate LDR with higher and lower exposure levels.
-Then the two exposure levels are combined to generate the HDR image.
+Then the two exposure levels are combined to generate the HDR.
+Wang et al.
+[94] proposed an unsupervised approach for HDR reconstruction.
+They train the model in such a way that when the HDR is re-projected to LDR using a camera response function model it becomes indistinguishable from the original LDR ( Sec.
+S1 for related works on non-learning methods).
 GAN-Based Approaches.
-Generative Adversarial Networks (GAN) [21, 3] architectures are well-explored in this domain.
+Generative Adversarial Networks [26, 2, 94] are well-explored in this domain.
 Niu et al.
@@
 Li et al.
-[49] proposed a GAN-based method for unpaired multi-exposure LDR to HDR translation.
-The method introduces modified GAN loss and novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR.
+[61] proposed a GAN-based unsupervised method for unpaired multi-exposure LDR-to-HDR translation.
+The method introduces modified GAN loss and a novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR.
 Wu et al.
-[82] proposed CycleGAN-like architecture for low-light image enhancement task using unpaired data for training.
-The method fuses the Retinex theory with CycleGAN concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR image.
+[102] proposed CycleGAN-like architecture for low-light image enhancement tasks using unpaired data for training.
+The method fuses the Retinex theory [54] with CycleGAN [117] concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR.
+Semantic and Knowledge-Based Approaches.
+Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction.
+Wang et al.
+[96] proposed a method that approximates the inverse of the camera pipeline.
+Their knowledge-inspired block uses the knowledge of image formation to address three tasks during HDR reconstruction: missing details recovery, adjustment of image parameters, and reduction of image noise.
+Goswami et al.
+[28] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera.
+The proposed method works in two stages: first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions.
+Liu et al.
+[65] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR.
+The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
+Limitations.
+Our analysis reveals that the methods based on single-exposed LDR fail to preserve image characteristics such as the levels of color hue and saturation, contrast, and brightness intensity in HDR.
+On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts such as halo in edges and boundaries, outlier pixel intensities, unwanted repeated pattens, missing details in shadow and highly bright regions, irregular color and texture transitions, and ghosting effects from dynamic scenes.
+While most methods produce excellent visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process that could significantly improve the quality of HDR output.
+Finally, LLM- and VLM-based [114, 105] perception capabilities have not been utilized in this problem area, which might improve performance.
+Tab.
+1 summarizes a comparison between the state-of-the-art and our method in terms of various parameters.
 Table 1: Summary of recent state-of-the-art methods.
-I/O: LDR used as input (SE: Single-exposed and ME: Multi-exposed), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), Unpaired: Uses unpaired data, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts, TM: Also performs tone-mapping i.e.
+I/O: LDR used as input, single-exposed (SE) and multi-exposed (ME), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), UP: Can be trained with unpaired data, LLM: Uses LLM-based perception, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts in heavily over/underexposed areas, TM: Also performs tone-mapping i.e.
 HDR $\rightarrow$ LDR.
-Method I/O O/P Unpaired Context Semantics Artifacts TM
+Method I/O O/P UP LLM Context Semantics Artifacts TM
 PSENet [76] SE D
@@
 Ghost-free HDR [65] ME D
+GlowGAN-ITM [94] SE D
 DITMO [28] SE I
-Ours SE D
-Semantic and Knowledge-Based Approaches.
-Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction.
-Wang et al.
-[77] proposed a method which approximates the inverse of the camera pipeline.
-Their knowledge-inspired block uses the image formation knowledge to address three tasks during HDR reconstruction: missing details recovery, image parameters adjustment, and image noise reduction.
-Goswami et al.
-[23] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera.
-The proposed method works in two stages - first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions.
-Liu et al.
-[52] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR.
-The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
-Limitations.
-Our analysis reveals that methods based on single-exposed LDR fail to preserve image characteristics such as color hue and saturation, color contrast, and intensity brightness levels in the HDR.
-On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts, missing details and ghosting effects.
-While most methods produce excellent results in terms of visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images.
-Table 1 summarizes a comparison between recent state-of-the-art and our method in terms of various parameters.
-Figure 1: Overview of the proposed method architecture (cf., Section 3.1) where $x$ and $y$ represent LDR and HDR images, respectively.
-The method is trained with five objectives: adversarial, cycle consistency, identity, contrastive, and semantic segmentation (cf., Section 3.2).
+LLM-HDR (Ours) SE D
+Figure 2: Overview of the proposed method architecture (Sec.
+3.1, Sec.
+S2.2) where $x$ and $y$ represent LDR and HDR images, respectively.
+The method is trained with six objectives: adversarial, cycle consistency, identity, LLM-based, contrastive, and semantic segmentation (Sec.
+3.2, Sec.
+S2.3).
 3 Method
@@
 3.1 Architecture Modules
-We adopt a CycleGAN [89] architecture as the basis of our method depicted in Figure 1.
-The network includes two generators and two discriminators (cf., Supplementary for brief background on cycle consistency).
+We adopt an cycle-consistent adversarial [117] architecture as the basis of our method depicted in Fig.
+2.
+The network includes two generators and two discriminators.
+The network also takes advantage of CLIP encoders and LLM-based perception to further bridge the gap between semantic and perceptual differences between input and generated images (Sec.
+S2.1 for a concise background on the key concepts employed in our method).
 Let us denote the two domains as $X$ for LDR images and $Y$ for HDR images.
@@
 Hence, the feedback block not only iterates over its own output but also re-runs the decoder situated ahead of it in each iteration while keeping the encoder frozen until the feedback iteration completes.
-The feedback is implemented with a ConvLSTM [30, 50] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (cf., Supplementary for visual depiction of the proposed generators and implementation details).
-Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [49, 32] (cf., Supplementary for implementation details).
+The feedback is implemented with a ConvLSTM [36, 63] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (Sec.
+S2.2 for visual depiction of the proposed generators and implementation details).
+Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [61, 39] (Sec. S2.2 for implementation details).
+LLM.
+We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each iteration of the training process.
+We prompt the LLM to provide the pixels identified as artifacts, and over/underexposed regions for both the reconstructed HDR and LDR.
+Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the pixels proposed by the LLM are used in a LLM-based loss function.
+In addition, the LLM generates a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps.
+The artifact map is multiplied with the input features to the feedback mechanism of the generators $G_{Y}$ and $G_{X}$.
+The underexposed map is fed into each of the skip connections (levels $1$ to $3$) in the generator $G_{Y}$ only and the overexposed map is multiplied with the input features to the bottleneck layer of this generator.
+We use element-wise multiplication as a gating mechanism to emphasize the selected areas in the maps.
+The rationale behind the separated saliency map fusion is based on the proposed U-Net architecture where distorted pixels are mostly refined by the feedback mechanism and later layers, high-intensity lighting details are generally recovered by the middle or later layers, and low-intensity shadows are reconstructed realistically at the starting and ending layers (Sec.
+S2.2 for visual depiction of the proposed LLM module and implementation details).
 Encoders.
 We also introduce a CLIP embeddings encoder $E$.
-Specifically we use the pre-trained Vision Transformer (ViT) [2] CLIP encoder from [66] which helps in extracting image embeddings with both local and global semantic context.
-For the forward cycle consistency we use $E(x)$ and $E(y)$.
-Then we add the embeddings from the encoder and feed them back to the bottleneck layer of the $G_{Y}$ decoder.
-Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(x)$ and feed them back to the bottleneck layer of the $G_{X}$ decoder.
+Specifically, we use a pre-trained CLIP encoder [81] to extract image embeddings with both local and global semantic context.
+For the forward cycle consistency we use $E(x)$ and $E(\hat{y})$.
+Then we add the embeddings and feed them back to the bottleneck layer of the $G_{Y}$ decoder.
+Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(\hat{x})$ and feed them back to the bottleneck layer of the $G_{X}$ decoder (Sec.
+S2.2 for implementation details).
 3.2 Loss Functions
-We use five loss functions to train the proposed method.
-Two are standard CycleGAN loss functions, i.e., adversarial loss [20] and cycle consistency loss [89].
+We introduce six loss functions to train the proposed method.
+We introduce three novel loss functions for HDR reconstruction - LLM-based, contrastive, and semantic segmentation loss.
+Two are standard loss functions, i.e., adversarial loss [25] and cycle consistency loss [117].
 Inspired from [117] we also use an identity loss [93].
-Finally, we introduce two novel loss functions for HDR reconstruction, i.e., contrastive and semantic segmentation loss.
-The loss is calculated on tone-mapped versions of the reconstructed and real HDR images.
-The tone-mapping is performed on the basis of the $\mu$-law [34].
-This is done to avoid the high intensity pixels of HDR images that can distort the loss calculation.
+The calculations in all loss functions are based on tone-mapped versions of the reconstructed and real HDR images.
+This tone-mapping is performed on the basis of the $\mu$-law [43] and is done to avoid the high-intensity pixels of HDR images that can distort the loss calculation.
 The tone-mapping operator $T$ can be represented as follows:
@@
 where the amount of compression $\mu$ is 5000 following [47].
+LLM Loss.
+Given a reconstructed HDR/LDR image, this loss is based on pixels identified by the LLM as artifacts and over/underexposed regions during Q&A sessions (using human-defined knowledge).
+The loss is defined as:
+$\mathcal{L}_{\text{llm}}=3\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+2\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+1.5\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
+where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed.
+The weights on each of the three components are selected after thorough ablations as well as theoretical considerations.
+Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye.
+Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component.
+(Sec.
+S2.3 for visual depiction of the LLM loss).
+This loss is calculated in the backward cycle with $\hat{x}$ but only with the artifact component.
+The rationale behind this decision lies in the nature of the output, i.e., LDR can be overexposed or underexposed.
+Contrastive Loss.
+This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains.
+Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [10].
+Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels.
+This is done by spreading out the frequent pixel intensity values across $256$ bins.
+Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image.
+For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as follows:
+$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (3)
+[…]
+$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (4)
+where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs.
+This loss is calculated in both the forward and backward cycles with the input and output being swapped.
+This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (Secs.
+S2.3 and S3.3 for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
+Semantic Segmentation Loss.
+This loss is based on segmentation masks.
+The Mean Intersection over Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks.
+Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [10].
+We choose equalized images instead of original LDR because segmentation of low-light or extremely bright images does not yield good results.
+We use Segment Anything (SAM) [50] to generate segmentation classes for the histogram-equalized LDR and reconstructed tone-mapped HDR images (Sec.
+S2.3 for visual depiction of the semantic segmentation loss).
+This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images.
+We can define the IoU metric as:
+$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (5)
+where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively.
+$\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$.
+The mean IoU over all segmentation classes can be formulated as:
+$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (6)
+We define the semantic segmentation loss as:
+$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (7)
+This loss is calculated in both forward and backward cycles with the input and output being swapped (Sec. S2.3 for visual depiction of the semantic segmentation loss).
 Adversarial Loss. We apply adversarial loss to both mappings. The mapping from LDR to HDR domain, i.e., $G_{Y}:X\rightarrow Y$ with the discriminator $D_{Y}$, can be expressed as:
@@
 Cycle Consistency Loss.
-The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each another.
+The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each other.
 Hence, we also incorporate a cycle consistency loss to prevent mutual contradiction of the learned mappings $G_{Y}$ and $G_{X}$.
@@
 $\begin{split}\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})=\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(G_{Y}(x))-x\|_{1}\right]+\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(G_{X}(y))-y\|_{1}\right],\end{split}$ (10)
-where $\|\cdot\|_{1}$ is the L1 norm (cf., Supplementary for visual depiction of the cycle consistency loss).
+where $\|\cdot\|_{1}$ is the L1 norm (Sec. S2.3 for visual depiction of the cycle consistency loss).
 Identity Loss.
-For LDR $\leftrightarrow$ HDR translation tasks, we also find that adversarial and cycle consistency formulations alone cannot preserve the color and hue information.
+For LDR $\leftrightarrow$ HDR translation tasks, we also found that adversarial and cycle consistency formulations alone cannot preserve the color and hue information.
 This is due to incorrect mapping of color shades from LDR to HDR domains by the generators stemming from the underlying difference in dynamic ranges.
@@
 This loss can be expressed as:
-$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (5)
-Contrastive Loss.
-This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains.
-Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [8].
-Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels.
-This is done by spreading out the frequent pixel intensity values across $256$ bins.
-Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image.
-For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(y)$ we first define the cosine similarity between them as follows:
-$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (6)
-[…]
-$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (7)
-where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs.
-This loss is calculated in both the forward and backward cycles with the input and output being swapped.
-This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (cf., Supplementary for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
-Semantic Segmentation Loss.
-This loss is based on segmentation masks.
-The Mean Intersection of Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks.
-Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [8].
-We choose equalized images instead of original LDR because segmentation on low-light or extreme brightness images does not yield good results.
-We use Segment Anything (SAM) [42] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images (cf., Supplementary for visual depiction of the semantic segmentation loss).
-This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images.
-We can define the IoU metric as:
-$\text{IoU}_{c}=\frac{x_{c}\cap y_{c}}{x_{c}\cup y_{c}},$ (8)
-where $x_{c}$ and $y_{c}$ represent the segmentations for class $c$ in image $x$ and $y$, respectively.
-$\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$.
-The mean IoU over all segmentation classes can be formulated as:
-$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (9)
-We represent the semantic segmentation loss as:
-$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (10)
-$\mathcal{L}_{\text{full}}=\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}$ (11)
+$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (11)
 Final Loss.
 The full objective is expressed in Eq.
-(11), where $\lambda$ scales the relative importance of the cycle consistency and identity loss.
+12, where $\lambda$ scales the relative importance of the cycle consistency and identity loss.
 $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [117].
 $\alpha$ and $\beta$ are weights for the contrastive and semantic segmentation loss, respectively.
-Both values are set to $2$ (cf., Supplementary for ablation experiments for $\alpha$ and $\beta$).
+Both values are set to $2$ (Sec.
+S3.3 for ablation experiments for $\alpha$ and $\beta$).
+$\begin{split}\mathcal{L}_{\text{full}}=\\ \mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\\ \lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\\ \alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}+\mathcal{L}_{\text{llm}}\end{split}$ (12)
 4 Experiments
-4.1 Implementation
-The proposed method is implemented in PyTorch [^1] https://pytorch.org/ on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD.
+We discuss the experimental setup details in Sec.
+S2.4.
+Datasets.
+For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [16], NTIRE [79], and HDR-Synth & HDR-Real [64] paired datasets.
+HDRTV has $1235$ samples for training and $117$ for testing.
+The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples.
+HDR-Synth & HDR-Real dataset consists of $20537$ samples.
+Other paired datasets used in our evaluations are DrTMO [23] ($1043$ samples), Kalantari [44] ($89$ samples), HDR-Eye [51] (46 samples), and LDR-HDR Pair [40] ($176$ samples).
+We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity.
+We used the pre-defined train/test sets of HDRTV and NTIRE.
+For HDR-Synth & HDR-Real, we performed a random $80/20$ split for training and testing in all experiments unless specified otherwise.
+For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR.
+For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [10] for datasets with only single-exposed LDR images.
+Metrics.
+We use four metrics to report the results.
+High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [68] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model.
+Structural Similarity Index Measure (SSIM) [99, 100, 101] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information.
+Peak Signal-to-Noise Ratio (PSNR) [31] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^1] Unless specified otherwise, for HDR evaluation PSNR is computed on tone-mapped original and reconstructed HDR image pairs..
+Learned Perceptual Image Patch Similarity (LPIPS) [115] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge-based contributions.
+This metric aligns with the perceptual judgement of humans.
+Table 2: HDR reconstruction result.
+Comparison with supervised (gray) and unsupervised, weakly-supervised and self-supervised (black) learning methods trained and evaluated on the paired datasets HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64].
+LP: Supervised (S), unsupervised (US), weakly-supervised (WS), and self-supervised (SS).
+[…]
+HDRCNN (ACM TOG’17[22] S 20.12 0.724 0.422 60.12
+DrTMO (ACM TOG’17[23] S 20.43 0.711 0.412 60.81
+ExpandNet (Eurographics’18[69] S 21.44 0.854 0.451 58.12
+FHDR (GlobalSIP’19[47] S 20.84 0.841 0.307 62.45
+SingleHDR (CVPR’20[64] S 21.01 0.851 0.402 64.21
+Two-stage (CVPRW’21[1] S 34.29 0.856 0.287 61.78
+HDR-GAN (IEEE TIP’21[77] S 30.11 0.912 0.411 65.01
+KUNet (IJCAI’22[96] S 37.21 0.974 0.051 62.14
+Ghost-free (ECCV’22[65] S 40.32 0.966 0.043 63.02
+ArtHDR-Net (APSIPA’23[5] S 37.12 0.961 0.321 63.43
+HistoHDR-Net (ICIP’24[6] S 38.21 0.968 0.301 65.15
+UPHDR-GAN (IEEE TCSVT’22[61] US 39.98 0.965 0.044 63.54
+PSENet (WACV’23[76] US 27.35 0.856 0.274 62.89
+SingleHDR(W) (WACV’23[55] WS 30.79 0.952 0.055 62.95
+GlowGAN-ITM (ICCV’23[94] US 30.19 0.901 0.064 60.05
+SelfHDR (ICLR’24[116] SS 39.51 0.972 0.037 64.77
+LLM-HDR (Ours) SS 40.11 0.979 0.020 68.51
+5 Results
+5.1 HDR Reconstruction
+For an extensive comparison we selected a combination of methods that utilize different learning paradigms including, HDRCNN [22], DrTMO [23], ExpandNet [69], FHDR [47], SingleHDR [64], Two-stage [1], KUNet [96], HistoHDR-Net [6], HDR-GAN [77], ArtHDR-Net [5], Ghost-free HDR [65], SelfHDR [116], UPHDR-GAN [61], GlowGAN-ITM [94], SingleHDR(W) [55], and PSENet [76].
+These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised, weakly-supervised, self-supervised, unpaired learning, and knowledge-conditioned.
+Paired Datasets.
+Tab.
+2 summarizes the results of evaluation on HDRTV [16], NTIRE [79], HDR-Synth & HDR-Real [64] paired datasets, the most widely used datasets by the state-of-the-art (Sec.
+S3.1 for an extensive qualitative evaluation).
+In the top part of the table, we compare our method with $11$ supervised learning approaches.
+The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2.
+In terms of PSNR, it is second best after Ghost-free HDR.
+In the bottom part of the table, we compare our method with $5$ unsupervised, weakly-supervised, and self-supervised learning methods.
+The proposed method outperforms those approaches on all evaluation metrics.
+It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training.
+However, in all experiments, our method is trained in a self-supervised fashion with unpaired samples (Sec.
+3).
+Mixed Datasets.
+To further demonstrate the strength of our approach we extended the train set used in the previous experiment (paired data) with: 1) The HDR images from DrTMO [23], Kalantari [44], HDR-Eye [51], LDR-HDR Pair [40], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images.
+There is no overlap between LDR and HDR in this additional data, i.e., it is an unpaired dataset.
+The final dataset used for training in this experiment has an approximately 87% overlap between LDR and HDR images.
+We kept the test set identical to the previous experiment, reported in Tab.
+2, in order to compare with the state-of-the-art methods based paired data.
+Given that the only other methods that support this type of unpaired train data are UPHDR-GAN [61] and GlowGAN-ITM [94], we perform a direct comparison with these approaches in Tab.
+3.
+The results demonstrate that our method outperforms all unpaired and paired data methods, setting new state-of-the-art on all metrics (Sec.
+S3.1 for more mixed data experiments).
+Table 3: HDR reconstruction results. Comparison with unsupervised learning methods trained on mixed datasets.
+[…]
+UPHDR-GAN [61] 42.51 0.980 0.043 66.13
+GlowGAN-ITM [94] 33.45 0.910 0.061 62.13
+LLM-HDR (Ours) 43.96 0.989 0.020 69.32
+Table 4: HDR reconstruction results. Comparison with unsupervised learning methods trained on unpaired datasets.
+[…]
+UPHDR-GAN [61] 37.51 0.941 0.101 61.63
+GlowGAN-ITM [94] 32.15 0.923 0.137 60.11
+LLM-HDR (Ours) 39.32 0.952 0.082 64.32
+Unpaired Datasets.
+We performed cross-dataset experiment where LDR and HDR for training are sourced from different datasets.
+We kept the test set identical to the previous experiments.
+LDR images are sourced from HDRTV [16] and NTIRE [79], while the HDR are taken from HDR-Synth & HDR-Real [64].
+The direct comparison with UPHDR-GAN [61] and GlowGAN-ITM [94] which have unpaired training capabilities is reported in Tab.
+4.
+The results demonstrate that our method outperforms both unpaired data methods on all metrics.
+5.2 LDR Reconstruction
+We report results for LDR reconstruction in Tab.
+5 and compare with the state-of-the-art tone-mapping operators $\mu$-law [43], Reinhard’s [84], and Photomatix [34].
+We also consider the HDR-to-LDR prior from GlowGAN [94] which is used to re-project reconstructed HDR to LDR space for inverse learning.
+The test data is identical to the one used in the HDR reconstruction experiments.
+The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS.
+In terms of PSNR, it is second best after Photomatix (Sec.
+S3.2 for more LDR reconstruction experiments).
+Table 5: LDR reconstruction results. Comparison with the state-of-the-art tone-mapping operators and GlowGAN-prior [94].
+[…]
+$\mu$-law operator [43] 28.68 0.891 0.311
+Reinhard’s operator [84] 30.12 0.903 0.342
+Photomatix [34] 32.11 0.922 0.136
+GlowGAN-prior [94] 32.03 0.936 0.108
+LLM-HDR (Ours) 31.81 0.942 0.091
+5.3 Ablation Results
+Generators.
+We propose a novel generator that modifies the U-Net architecture by introducing an artifact-aware feedback mechanism (Sec.
+3.1).
+The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult.
+To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [55] method with our U-Net generator.
+We see that this leads to improvements in all but one metrics as shown in Tab.
+6.
+The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W).
+Fig.
+3 illustrates the improvement in SingleHDR(W) [55] when we use the proposed feedback U-Net instead of the original U-Net of SingleHDR(W).
+The original U-Net produces many artifacts in the output HDR images whereas our modified version with feedback reconstructs artifact-free HDR images.
+Sec.
+S3.3 provides extensive ablation experiments for the proposed architecture and loss functions.
+Figure 3: Comparison of the SingleHDR(W) [55] U-Net with and without our feedback mechanism on images from the DrTMO [23].
+6 Conclusion
+[…]
+Table 6: Quantitative evaluation of our feedback U-Net-based generators for HDR reconstruction.
+For evaluation we use the dataset employed by SingleHDR(W), i.e., DrTMO [23].
+The models marked with “(mod)” use feedback U-Net-based generators and the rest use the same U-Net generator as SingleHDR(W).
+[…]
+SingleHDR(W) [55] 30.79 0.952 0.055 62.95
+SingleHDR(W) (mod) 30.03 0.961 0.036 63.91
+LLM-HDR 35.11 0.969 0.029 67.88
+LLM-HDR (mod) 43.01 0.988 0.023 68.21
+Supplementary Material
+S1 Related Work
+This section complements Sec.
+2 in the main paper and includes an overview of non-learning approaches for HDR reconstruction.
+The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [42].
+This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image.
+Another approach is using histogram equalization [20] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image.
+Gradient domain manipulation [52] techniques enhance the granular details of an LDR image and expand its dynamic range.
+Pyramid-based image fusion [46] is a technique where image features are extracted into Laplacian pyramids [46] and then fused for the HDR.
+Some methods use Retinex-based [48, 71] tuning to effectively produce and display HDR images on consumer screens [48, 71].
+Intensity mapping function-based approaches [110, 21, 80] use transformations on image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels.
+The methods in this category, although time-efficient, face some issues including ghosting effects, halo artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
+S2 Method
+This section complements Sec.
+3 in the main paper.
+It provides a concise background on the key concepts employed in our method, i.e., cycle consistency [117], Large Language Models [75, 105], contrastive language-image pretraining [81, 62, 112], and image semantic segmentation [50].
+The section also offers implementation details for the architecture modules.
+S2.1 Background
+Cycle Consistency.
+The concept of cycle consistency in image-to-image [117] and video-to-video [17] translation tasks is a powerful and elegant design for unpaired data.
+Previously, cycle consistency has been applied to paired LDR $\leftrightarrow$ HDR translation [106], where the authors used $3$ LDR images with different exposures as input to reconstruct an HDR.
+The reconstructed HDR is then used by $3$ different generators to reconstruct the original $3$ multi-exposed LDR images.
+However, the focus of our work is on the more general task of unpaired LDR $\leftrightarrow$ HDR translation.
+Given the images $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$, the goal is to design two generators, i.e., $G_{Y}$ and $G_{X}$ such that, $G_{Y}(x)$rightarrow $y$ and $G_{X}(y)$rightarrow $x$.
+Fig.
+S1 depicts this concept for the LDR-to-HDR translation task.
+Figure S1: $G_{Y}$ translates LDR to HDR images and $G_{X}$ HDR to LDR. The LDR images are from the LDR-HDR pair [40] dataset.
+LLM and VLM.
+Large Language Models [75] and Vision-Language Models [114, 105] are trained on vast amounts of textual and/or visual data (i.e., images and videos) and achieve excellent zero-shot performance.
+We leverage the text and visual perception capabilities of an LLM to evaluate the reconstruction of our method and refine it with better awareness of artifacts and over/underexposed regions.
+Contrastive Language-Image Pretraining.
+CLIP [81] is a method that finds the closeness between visual and textual data in the embedding space.
+The method is trained on image-text pairs to associate images with their descriptions.
+The CLIP loss maximizes the cosine similarity between correct (positive) pairs and minimize the similarity between incorrect (negative) pairs.
+We extract the CLIP embeddings from both the LDR and reconstructed HDR image pairs and maximize their cosine similarity in the embedding space for preservation of semantic information across domains.
+We also combine the CLIP embeddings extracted from the LDR and reconstructed HDR to guide the decoder of the generators in the next step of the reconstruction.
+Semantic Segmentation.
+Semantic segmentation [50, 83, 59] is a vision problem where each pixel of an image or video frame is segmented into meaningful sub-parts such as objects, roads, water bodies, buildings, humans, and animals.
+Segment Anything (SAM) [50] is a state-of-the-art method for semantic segmentation of images.
+Segmented objects in a scene not only inform the number and location of semantically meaningful concepts, but also provide a holistic semantic description of the scene.
+Therefore, we can use that information to improve image reconstruction.
+We introduce a loss on the basis of the Mean Intersection over Union metric to preserve the image integrity (based on the segmented objects) between the two domain $X$ and $Y$.
+Figure S2: Overview of the proposed generators based on our novel feedback based U-Net architecture.
+Left part (blue) is the encoder and right part (red) is the decoder.
+The decoder is inside the feedback iteration loop.
+Figure S3: Overview of the proposed LLM module.
+Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not.
+Must response with 1) Yes or No only, 2) If Yes, can you get me the saliency maps of the artifacts (in red), overexposed (in blue) and underexposed (in yellow) areas of this image?”, System: “Outputs the saliency maps for the detected artifacts in red, overexposed pixels in blue, and underexposed pixels in yellow.”
+S2.2 Architecture Modules
+Generators.
+The U-Net of each generator consists of $7$ levels.
+Each level consists of $2$ Conv layers with $3\times 3$ kernel followed by a ReLU activation and a BN layer.
+The input images of size $512\times 512$ are first converted into $32$-dimensional features.
+The features are doubled in each level of the encoder until reaching size of $512$.
+The decoder levels then up-sample the data by performing channel-wise concatenation with the extracted feature from previous levels of the encoder.
+The intuition behind using feedback mechanism is to refine the decoder layers of the generator based on the input from encoder layers and the feedback mechanism itself.
+The output of each iteration of the feedback mechanism is not only fed into the first level of the decoder block, but also sent back through the hidden state to guide its next input.
+Feedback mechanisms perform better than simple feed-forward networks with less trainable parameters [63].
+The feedback network consists of $3$ dilated dense blocks each with dilation rate of $3$.
+The dilated blocks are made of $3\times 3$ Conv layers followed by ReLU activation function.
+Each dilated block contains two $1\times 1$ feature compression layers at the start and end.
+Dilation enhances the network’s receptive field which helps to capture broader context from the input features without increasing the computational complexity.
+Finally, the feedback starts with a $1\times 1$ Conv and ends with $3\times 3$ Conv for compression and final processing, respectively.
+Fig.
+S2 depicts and overview of the proposed generators architecture.
+Figure S4: Depiction of the LLM-based loss $\mathcal{L}_{\text{llm}}$.
+Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not.
+Must response with 1) Yes or No only, 2) If Yes, return the [artifact areas number of pixels, total number of pixels].
+Also highlight the artifact areas with pink.
+Similarly find if there are any over/under exposure areas in the scene and return the [pixels in overexposure, pixels in underexposure areas] and highlight the overexposed areas in blue and underexposed areas in green.”, System: “Outputs the number of pixels with the pixel list (in .txt files) for the detected artifacts in pink, overexposed pixels in blue, and underexposed pixels in green.”
+Figure S5: Depiction of the contrastive loss $\mathcal{L}_{\text{con}}$.
+Positive (green) and negative (red) pairs in a batch.
+We use a histogram-equalized version of the LDR processed using the OpenCV function equalizeHist [10].
+Figure S6: Depiction of the semantic segmentation loss $\mathcal{L}_{\text{sem}}$.
+We use Segment Anything (SAM) [50] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images.
+Figure S7: Depiction of the cycle consistency loss $\mathcal{L}_{\text{cyc}}$ using an image from the DrTMO [23] dataset.
+Discriminators.
+The discriminators consist of $5$ Conv layers.
+We input the reconstructed LDR/HDR along with a real (unpaired) LDR/HDR in the respective discriminator.
+The Conv layers have $4\times 4$ filters and stride of $2$ for the first $3$ layers and $1$ henceforth.
+The output of the first layer has features of size $64$, which is doubled in each of the next layers until $512$.
+The output of the fifth layer is of size $1$.
+The first four layers have LeakyReLU activation with the final layer using a Sigmoid activation for outputting the probabilities of a particular image being real or fake.
+The training process is stabilized using BN following all the Conv layers except the first one.
+LLM.
+The artifact-aware saliency map $LLM_{\text{af}}$ is extracted as feature of size $512$ with $1\times 1$ Conv and we perform an element-wise multiplication with the output features of the bottleneck layer (i.e., layer $4$) of the U-Net generators $G_{Y}$ and $G_{X}$ which goes into the feedback mechanism.
+The exposure-aware saliency map $LLM_{\text{ox}}$ (for overexposure) (which goes into the generator $G_{Y}$ only) is extracted as feature of size $256$ with $1\times 1$ Conv and fused with the output features of layer $3$ henceforth concatenated with the encoder embeddings which goes into the bottleneck layer.
+Finally, $LLM_{\text{ux}}$ (for underexposure) is extracted as feature of sizes $64$, $128$, and $256$, with $1\times 1$ Conv and fused with the skip connections in each level of the U-Net generator $G_{Y}$ (i.e., level 1, 2, and 3).
+Fig.
+S3 depicts the overview of the proposed $LLM$ module.
+Encoders.
+The encoders extract embeddings of size $512$ from the input and output domain images.
+Given the input size of the generators, we re-project these embeddings to a size of $256$ using $1\times 1$ Conv.
+Then we add the embeddings and concatenate them with the input features of the bottleneck layer (i.e., layer $4$) of our feedback-based U-Net generators.
+S2.3 Loss Functions
+Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM-based loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
+S2.4 Implementation Details
+The proposed method is implemented in PyTorch [78] on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD.
 We used Adam optimizer [49] to train the models for $170$ epochs.
@@
 The HDR images used in the text are tone-mapped with Reinhard’s operator [84].
-4.2 Datasets
-For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [13], NTIRE [64], and HDR-Synth & HDR-Real [51] paired datasets.
-HDRTV has $1235$ samples for training and $117$ for testing.
-The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples.
-HDR-Synth & HDR-Real dataset consists of $20537$ samples.
-Other paired datasets used in our evaluations are DrTMO [19] ($1043$ samples), Kalantari [36] ($89$ samples), HDR-Eye [43] (46 samples), and LDR-HDR Pair [33] ($176$ samples).
-We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity.
-We perform $80/20$ split for training and testing unless specified otherwise.
-For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR.
-For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [8] for datasets with only single-exposed LDR images.
-4.3 Metrics
-We use four metrics to report the results.
-High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [55] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model.
-Structural Similarity Index Measure (SSIM) [79, 80, 81] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information.
-Peak Signal-to-Noise Ratio (PSNR) [26] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^2] PSNR is computed on $\mu$-law based tone-mapped HDR images..
-Learned Perceptual Image Patch Similarity (LPIPS) [87] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge based contributions.
-This metric aligns with the perceptual judgement of humans.
-Table 2: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach.
-Quantitative comparison with supervised (gray) and unsupervised/weakly-supervised/self-supervised (black) learning methods trained on the paired datasets HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51].
+S3 Results
+This section complements Sec.
+5 in the main paper.
+It provides additional quantitative evaluation of our method for both HDR and LDR reconstruction.
+The section also includes an extensive qualitative and ablation results.
+S3.1 HDR Reconstruction
+Tab.
+S1 provides the results from an additional mixed data experiment.
+The train set used in this experiment is: 1) The HDR images from DrTMO [23], Kalantari [44], HDR-Eye [51], LDR-HDR Pair [40], and GTA-HDR [7] ($1500$ random samples), i.e., $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., $2677$ LDR images.
+In this dataset, a total of $1177$LDR,HDR} pairs overlap.
+Given that the only other methods that support this type of unpaired train data are UPHDR-GAN [61] and GlowGAN-ITM [94], we perform a direct comparison with these two approaches.
+The results demonstrate that our method again outperforms UPHDR-GAN and GlowGAN-ITM on all metrics.
+Table S1: HDR reconstruction results. Comparison with unsupervised learning methods trained on mixed datasets.
 […]
-HDRCNN [18] 20.12 0.724 0.422 60.12
-DrTMO [19] 20.43 0.711 0.412 60.81
-ExpandNet [56] 21.44 0.854 0.451 58.12
-FHDR [39] 20.84 0.841 0.307 62.45
-SingleHDR [51] 21.01 0.851 0.402 64.21
-Two-stage [1] 34.29 0.856 0.287 61.78
-HDR-GAN [63] 30.11 0.912 0.411 65.01
-KUNet [77] 37.21 0.974 0.051 62.14
-Ghost-free [52] 40.32 0.966 0.043 63.02
-ArtHDR-Net [6] 37.12 0.961 0.321 63.43
-HistoHDR-Net [5] 38.21 0.968 0.301 65.15
-UPHDR-GAN [49] 39.98 0.965 0.044 63.54
-PSENet [62] 27.35 0.856 0.274 62.89
-SingleHDR(W) [45] 30.79 0.952 0.055 62.95
-SelfHDR [88] 39.51 0.972 0.037 64.77
-Ours 40.01 0.979 0.021 68.41
-Table 3: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with UPHDR-GAN trained on unpaired datasets.
-[…]
-UPHDR-GAN [49] 42.51 0.980 0.043 66.13
-Ours 43.92 0.989 0.020 69.32
-5 Results
-5.1 HDR Reconstruction
-For an extensive comparison we choose a combination of methods that utilize different learning paradigms including, HDRCNN [18], DrTMO [19], ExpandNet [56], FHDR [39], SingleHDR [51], Two-stage [1], KUNet [77], HistoHDR-Net [5], HDR-GAN [63], ArtHDR-Net [6], Ghost-free HDR [52], SelfHDR [88], UPHDR-GAN [49], SingleHDR(W) [45], and PSENet [62].
-These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised/weakly-supervised/self-supervised, unpaired learning, and knowledge-conditioned.
-Table 4: LDR reconstruction (tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with the state-of-the-art one-mapping operators.
+UPHDR-GAN [61] 41.98 0.975 0.046 65.54
+GlowGAN-ITM [94] 32.18 0.905 0.066 61.89
+LLM-HDR (Ours) 42.01 0.988 0.023 68.81
+Figs.
+S12, S13, S14, S15, S16, S17 and S18 illustrate the quality of the HDR images reconstructed with our method.
+Our method mitigates artifacts such that the resulting HDR closely resembles the ground truth HDR.
+The finer textures, color hues, and shades are well preserved compared to the other methods.
+The granular level details in the bright and dark regions are also well reconstructed.
+S3.2 LDR Reconstruction
+Tab.
+S2 provides the results from an additional LDR reconstruction experiment.
+In this case, our method is trained with the data specified in the previous section.
+Similar to the LDR reconstruction experiment in the main paper, we compare the model with the state-of-the-art tone-mapping operators $\mu$-law [43], Reinhard’s [84], Photomatix [34], and GlowGAN-prior [94].
+The results again demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS.
+In terms of PSNR, it is third best after Photomatix and GlowGAN-prior.
+Table S2: LDR reconstruction results. Comparison with the state-of-the-art tone-mapping operators and GlowGAN-prior [94].
 […]
@@
 Photomatix [34] 32.11 0.922 0.136
-Ours 31.71 0.941 0.092
-Table 2 summarizes the results of evaluation on HDRTV [13], NTIRE [64], HDR-Synth & HDR-Real [51] paired datasets, the most widely used datasets by the state-of-the-art (cf., Supplementary for an extensive qualitative evaluation).
-In the top part of the table, we compare our method with $11$ supervised learning approaches.
-The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2.
-In terms of PSNR, it is second best after Ghost-free HDR.
-In the bottom part of the table, we compare our method with $4$ unsupervised/weakly-supervised/self-supervised learning methods.
-The proposed method outperforms those approaches on all evaluation metrics (cf., Supplementary for visual depiction of the quantitative HDR reconstruction results).
-It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training.
-However, in all experiments, our method is trained in a self-supervised fashion using only unpaired samples (cf., Section 3).
-To further demonstrate the strength of our approach we extended the train set used in the previous experiment with: 1) The HDR images from DrTMO [19], Kalantari [36], HDR-Eye [43], LDR-HDR Pair [33], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images.
-We kept the test set identical to the previous experiment in order to compare with the state-of-the-art paired data based methods.
-Given that the only other method that supports this type of unpaired train data is UPHDR-GAN, we perform a direct comparison with that approach in Table 3.
-The results demonstrate that our method outperforms UPHDR-GAN and sets new state-of-the-art on all metrics (cf., Supplementary for more unpaired data experiments).
-Table 5: Quantitative evaluation of our feedback U-Net-based generators for HDR reconstruction.
-For evaluation we use the dataset employed by SingleHDR(W), i.e., DrTMO [19].
-The models marked with “(mod)” use feedback U-Net-based generators and the rest use the same U-Net generator as SingleHDR(W).
-[…]
-SingleHDR(W) [45] 30.79 0.952 0.055 62.95
-SingleHDR(W) (mod) 30.03 0.961 0.036 63.91
-Ours 35.11 0.969 0.029 67.88
-Ours (mod) 43.01 0.988 0.023 68.21
-Table 6: Architecture ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
-Uo: The U-Net generator used in SingleHDR(W) [45], Uf: The proposed U-Net generator, E: CLIP embedding encoder.
+GlowGAN-prior [94] 31.91 0.931 0.112
+LLM-HDR (Ours) 31.11 0.939 0.095
+Table S3: Architecture ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
+Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware maps.
 […]
@@
 Uf 24.07 0.879 0.126 63.16
+Uf+LLM 26.12 0.880 0.121 64.32
 Uf+E 29.19 0.885 0.122 64.21
-Uf+E+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
-Uf+E+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
-Uo+E+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 35.15 0.953 0.042 67.18
-Ours 40.01 0.979 0.021 68.41
-5.2 LDR Reconstruction
-For a complete evaluation, we also report results for LDR image reconstruction in Table 4 and compare with the state-of-the-art tone-mapping operators $\mu$-law [34], Reinhard’s [68], and Photomatix [28].
-The test data is identical to the one used in the HDR reconstruction experiments.
-The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS.
-In terms of PSNR, it is second best after Photomatix [68] (cf., Supplementary for more LDR reconstruction experiments).
-5.3 Ablation Results
-Generators.
-We propose a novel generator that modifies the U-Net architecture by introducing a feedback mechanism (cf., Section 3.1).
-The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult.
-To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [45] method with our U-Net generator.
-We see that leads to improvement in all metrics as shown in Table 5.
-The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W) (cf., Supplementary for qualitative results in support of this quantitative evaluation).
+Uf+E+LLM 30.15 0.887 0.101 64.41
+Uf+E+LLM+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
+Uf+E+LLM+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
+Uf+E+LLM+$\mathcal{L}_{\text{llm}}$ 36.01 0.907 0.046 66.72
+Uf+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 40.01 0.979 0.021 68.41
+Uo+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 35.15 0.953 0.042 67.18
+Uf+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 40.11 0.979 0.020 68.51
+S3.3 Ablation Results
 Architecture.
 Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task.
-The most significant components are the CLIP embedding encoder and feedback U-Net generators (cf., Section 3.1).
+The most significant components are the LLM module, artifact-aware feedback and exposure-aware skip connection-based U-Net generators, and the CLIP embedding encoders (Sec.
+3.1).
 We perform a thorough ablation on each of the components to highlight the contributions.
-Table 6 summarizes the results where we add each component to the architecture as we go down.
+Tab.
+S3 summarizes the results where we add each component to the architecture as we go down.
 The first row lists the results from the original CycleGAN method.
-The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W) [45].
+The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W).
 The third row is the CycleGAN utilizing our feedback U-Net generator.
-This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2.
-The fourth row lists the results when we add the CLIP embedding encoder to the architecture.
-This results in an improvement in all metrics.
-The fifth and sixth rows depict the improvements after adding contrastive and semantic segmentation objectives in the architecture (cf., Section 3.2).
-The seventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W).
-Finally, the last row is the full architecture with all components which achieves the best performance on all metrics.
+This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2 (Q-Score).
+The fourth row shows the results when we add the LLM module to the architecture which shows some improvement in the PSNR, LPIPS, and Q-Score.
+The fifth row lists the results when we remove the LLM module and add the CLIP embedding encoders to the architecture.
+This results in an improvement in PSNR, SSIM, and Q-Score.
+The sixth row depicts good improvement in almost all the metrics as we add both CLIP encoders and LLM module to the architecture.
+The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM-based objectives (Sec.
+3.2) in the architecture, respectively.
+The tenth row shows the improvements when we add only the contrastive and semantic segmentation objectives in the architecture.
+The eleventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W).
+The last row is the full architecture with all components which achieves the best performance on all metrics.
 Loss Functions.
 We also study the loss functions and their contribution towards the overall results keeping the architectural components constant.
-Table 7 summarizes the results.
+Tab.
+S4 summarizes the results.
 The first row is the results from the standard adversarial and cycle consistency loss.
-In the second row we add the identity loss which improves all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{id}}$).
-In the third and forth rows we add the contrastive and semantic segmentation loss.
-This results in an improvement in all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$).
-Finally, adding all loss functions achieves the best results on all the metrics.
-Table 7: Loss functions ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
+In the second row we add the identity loss which improves all metrics.
+In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM-based objectives.
+This results in an improvement in all metrics.
+The sixth row shows the results when we add contrastive and semantic losses together which gives significant improvement in the metrics.
+Finally, adding all loss functions (seventh row) achieves the best results on all metrics.
+Fig.
+S9 visually validates the contribution of the identity loss $\mathcal{L}_{\text{id}}$ illustrating that it helps in recovering the hue and shades of original image in the reconstruction.
+Figs.
+S10 and S11 demonstrate similar results for the contrastive and semantic segmentation objectives.
+The contrastive loss $\mathcal{L}_{\text{con}}$ can recover high-level attributes such as color and texture information more accurately compared to our method without this loss component.
+We also observe that the semantic segmentation loss $\mathcal{L}_{\text{sem}}$ recovers the low-level attributes such as object boundaries and edges more vividly than our method without this loss component.
+These high/low-level attributes are of vital importance to downstream applications like object recognition.
+Table S4: Loss functions ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
 […]
 $\mathcal{L}_{\text{GAN}}$+$\mathcal{L}_{\text{cyc}}$ 24.55 0.878 0.125 62.11
-+$\mathcal{L}_{\text{id}}$ 29.19 0.885 0.122 64.21
++$\mathcal{L}_{\text{id}}$ 30.15 0.887 0.101 64.41
 +$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
 +$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
++$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{llm}}$ 36.01 0.907 0.046 66.72
 +$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 40.01 0.979 0.021 68.41
++$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 40.11 0.979 0.020 68.51
 We also perform separate tests on contrastive and semantic segmentation loss to examine the contribution of histogram-equalized LDR compared to original LDR images.
-We summarize these results in Tables 8 and 9.
+We summarize these results in Tabs.
+S6 and S5.
 In both cases the loss calculated between the histogram-equalized versions of the LDR (instead of original the LDR) and the recnstructed HDR is better.
 This is due to the fact that histogram-equalized LDR can reveal semantic information in extremely over/underexposed areas of the original LDR images which in turn improves the extracted CLIP embeddings for $\mathcal{L}_{\text{con}}$ and segmentation maps for $\mathcal{L}_{\text{sem}}$.
-Table 8: Contrastive loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
+Table S5: Semantic segmentation loss ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
+[…]
+LDR 38.89 0.956 0.033 66.92
+LDRHist 40.11 0.979 0.020 68.51
+Table S6: Contrastive loss ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
 […]
 LDR 39.12 0.966 0.030 67.11
-LDRHist 40.01 0.979 0.021 68.41
-Table 9: Segmentation loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
-[…]
-LDR 38.89 0.956 0.033 66.92
-LDRHist 40.01 0.979 0.021 68.41
-6 Conclusions
-Most current state-of-the-art methods for HDR reconstruction (inverse tone-mapping) require high-quality paired {LDR,HDR} datasets for model training.
-Consequently, the quality of the state-of-the-art in inverse tone-mapping depends on the quality of the available paired {LDR,HDR} datasets.
-Additionally, only a few approaches utilize semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images.
-[…]
+LDRHist 40.11 0.979 0.020 68.51
+Finally, we experiment with the weights for the loss components with respect to PSNR.
+We perform tests on the temperature parameter $\tau$ of the contrastive loss and the weight parameters $\alpha$ and $\beta$ for the loss components $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$.
+Fig.
+S8 shows the results.
+The temperature parameter $\tau$ gives the optimal PSNR with $\tau=0.08$ while the result degrades if we increase the value further.
+If we decrease the value, it gives a sub-optimal PSNR at $\tau=0.07$, so we fix the value to $\tau=0.08$ in all experiments.
+The values of $\alpha$ and $\beta$ give optimal PSNR for $\alpha=2$ and $\beta=2$.
+The weights for the LLM-based loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively.
+We conducted ablation experiments on these weights considering values from $1$ to $5$ and found that the selected combination of weights yield the best PSNR.
+Figure S8: Ablation results for the temperature parameter $\tau$ (a), and loss component weights $\alpha$ (b) and $\beta$ (c).
+S4 Limitations and Future Work
+Our method is designed for LDR $\leftrightarrow$ HDR bi-directional translation for images and can be further extended to video-based LDR $\leftrightarrow$ HDR translation with some modifications in the generators (e.g., using special ViT [32] or Timesformer [9]) models.
+Diffusion-based cycle consistency approach can also be explored for video-based LDR/HDR translation.
+Further research can also explore multi-task architectures along with dynamic range translation such as deblurring, denoising, super-resolution, and demosaicing tasks.
+The pre-processing step in our loss calculations for contrastive and semantic segmentation loss employs histogram equalization, which can be further investigated with methods like Gamma correction [58] or histogram matching [18].
+Another direction can be the integration of no-reference quality assessment [4] into the framework which will enable parallel training and prediction of the reconstructed LDR/HDR along with its quality score.
+More advanced LLM-based formulation such as explainability of the generation process and object’s material based specular reflections can also be introduced in the future.
+Moreover, the Differentiable Ray Tracing technology [24] used to generate realistic HDR images/videos in video games [7] to simulate the physics behind light, specular reflections and shadows can also be explored using deep learning concepts like Nvdiffrast [53] and Neural Radiance Fields (NeRFs) [73] in 2D domain.
+Figure S9: Ablation results for $\mathcal{L}_{\text{id}}$ with images from the HDR-Synth & HDR-Real dataset [64] and LDR-HDR pair datasets [40].
+Figure S10: Ablation results for $\mathcal{L}_{\text{con}}$ with images from the LDR-HDR pair [40] dataset.
+Figure S11: Ablation results for $\mathcal{L}_{\text{sem}}$ with images from the HDR-Synth & HDR-Real dataset [64] dataset.
+Figure S12: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S13: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S14: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S15: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S16: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S17: Examples of HDR images reconstructed with our method and recent state-of-the-art.
+Figure S18: Examples of HDR images reconstructed with our method and recent state-of-the-art.

### v3

@@
 In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing Contrastive Language-Image Pretraining (CLIP) [81] embeddings and loss based on semantic segmentation.
-Moreover, LLM-based artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation.
+Moreover, LLM-based reasoning of artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation.
 The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images.
@@
 […]
-  • We leverage LLM-based loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs. 3.1 and 3.2).
+  • We leverage LLM-based reasoning to obtain a loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs.
+3.1 and 3.2).
   • We propose a novel generator based on a modified U-Net architecture [86] that incorporates ConvLSTM-based artifact-aware feedback mechanism [36, 63] and exposure-aware skip connections to mitigate visual artifacts in the HDR reconstruction (Sec.
@@
 LLM.
-We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each iteration of the training process.
-We prompt the LLM to provide the pixels identified as artifacts, and over/underexposed regions for both the reconstructed HDR and LDR.
-Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the pixels proposed by the LLM are used in a LLM-based loss function.
-In addition, the LLM generates a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps.
+We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each epoch of the training process as early training outputs are noisy.
+Evaluating once in every epoch allows the LLM to assess relatively more stable and informative outputs reducing noise in its feedback.
+We prompt the LLM to provide the reasoning for whether the generated image has artifacts, and over/underexposed regions for both the reconstructed HDR and LDR.
+Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the reasoning of the images proposed by the LLM along with a heuristic based post-processing using OpenCV [10] to obtain the corresponding pixels are used in a LLM-based loss function.
+In addition, the LLM’s reasoning also gives way to the generation of a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps using the same heuristic based post-processing.
 The artifact map is multiplied with the input features to the feedback mechanism of the generators $G_{Y}$ and $G_{X}$.
@@
 where the amount of compression $\mu$ is 5000 following [47].
-LLM Loss.
-Given a reconstructed HDR/LDR image, this loss is based on pixels identified by the LLM as artifacts and over/underexposed regions during Q&A sessions (using human-defined knowledge).
+LLM Reasoning Inspired Loss.
+Given a reconstructed HDR/LDR image, this loss is based on pixels identified as artifacts and over/underexposed regions by the heuristic based post-processing using OpenCV [10] inspired by the LLM’s reasoning during Q&A sessions (using human-defined knowledge).
 The loss is defined as:
-$\mathcal{L}_{\text{llm}}=3\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+2\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+1.5\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
-where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed.
+$\mathcal{L}_{\text{llm}}=\delta_{1}\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+\delta_{2}\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+\delta_{3}\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
+where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed and generated by the heuristic based post-processing.
 The weights on each of the three components are selected after thorough ablations as well as theoretical considerations.
-Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye.
-Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component.
+Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye hence, $\delta_{1}$ is set to 3.
+Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component, i.e., $\delta_{2}$ and $\delta_{3}$ are set to 2 and 1.5 correspondingly.
 (Sec.
@@
 Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image.
-For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as follows:
-$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (3)
-[…]
-$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (4)
+For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as $\text{sim}(\bar{x},\bar{y})$ (Sec.
+S2.3 for full equation).
+We formulate the contrastive loss for an input batch as positive pairs of images (e.g., LDR and the corresponding reconstructed HDR) and negative pairs of images (e.g., each LDR with the rest of the LDR as well as the rest of the reconstructed HDR in the batch) as follows:
+$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (3)
 where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs.
@@
 This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images.
-We can define the IoU metric as:
-$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (5)
-where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively.
-$\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$.
-The mean IoU over all segmentation classes can be formulated as:
-$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (6)
+We can define the IoU metric as $\text{IoU}_{c}$.
+The mean IoU over all segmentation classes can be represented as mIoU (Sec.
+S2.3 for full equations).
 We define the semantic segmentation loss as:
@@
 The full objective is expressed in Eq.
-12, where $\lambda$ scales the relative importance of the cycle consistency and identity loss.
+9, where $\lambda$ scales the relative importance of the cycle consistency and identity loss.
 $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [117].
@@
 These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised, weakly-supervised, self-supervised, unpaired learning, and knowledge-conditioned.
+Figure 3: Examples of HDR images reconstructed with our method and recent state-of-the-art.
 Paired Datasets.
@@
 The proposed method outperforms those approaches on all evaluation metrics.
+Although the improvement is small for PSNR and SSIM but LPIPS and Q-Socre show considerable improvement, which is important to note as perceptual, semantic, and high dynamic qualities are captured by these two metrices only.
 It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training.
@@
 Figure S3: Overview of the proposed LLM module.
+Here LLM signifies the LLM based reasoning of the input image and the heuristic based post-processing with OpenCV [10] for generation of the saliency maps.
 Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not.
@@
 S2 depicts and overview of the proposed generators architecture.
-Figure S4: Depiction of the LLM-based loss $\mathcal{L}_{\text{llm}}$.
+Figure S4: Depiction of the LLM Reasoning Inspired loss $\mathcal{L}_{\text{llm}}$.
+Here LLM signifies the LLM based reasoning of the input image and the heuristic based post-processing with OpenCV [10] for generation of the affected pixel positions.
 Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not.
@@
 LLM.
+We use the reasoning from the LLM to determine whether there is any exposure/artifact issues in the generation.
+Henceforth, we use heuristic based OpenCV [10] functions to threshold the pixels to generate the exposure-aware masks.
+Moreover, we also use the Laplacian and GaussianBlur functions of OpenCV to generate the artifact-aware masks.
 The artifact-aware saliency map $LLM_{\text{af}}$ is extracted as feature of size $512$ with $1\times 1$ Conv and we perform an element-wise multiplication with the output features of the bottleneck layer (i.e., layer $4$) of the U-Net generators $G_{Y}$ and $G_{X}$ which goes into the feedback mechanism.
@@
 S2.3 Loss Functions
-Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM-based loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
+Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM Reasoning Inspired loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
+The $\text{IoU}_{c}$ can be formulated as:
+$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (10)
+where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively.
+$\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$.
+The $\text{mIoU}_{c}$ can be formulated as:
+$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (11)
+The cosine similarity $\text{sim}(\bar{x},\bar{y})$ can be formulated as:
+$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (12)
+[…]
 S2.4 Implementation Details
@@
 Figs.
-S12, S13, S14, S15, S16, S17 and S18 illustrate the quality of the HDR images reconstructed with our method.
+S12, S13, S14, S15, S16 and S17 illustrate the quality of the HDR images reconstructed with our method.
 Our method mitigates artifacts such that the resulting HDR closely resembles the ground truth HDR.
@@
 Table S3: Architecture ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
-Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware maps.
+Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware reasoning.
 […]
@@
 The sixth row depicts good improvement in almost all the metrics as we add both CLIP encoders and LLM module to the architecture.
-The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM-based objectives (Sec.
+The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM Reasoning Inspired objectives (Sec.
 3.2) in the architecture, respectively.
@@
 In the second row we add the identity loss which improves all metrics.
-In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM-based objectives.
+In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM Reasoning Inspired objectives.
 This results in an improvement in all metrics.
@@
 The values of $\alpha$ and $\beta$ give optimal PSNR for $\alpha=2$ and $\beta=2$.
-The weights for the LLM-based loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively.
+The weights for the LLM Reasoning Inspired loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively.
 We conducted ablation experiments on these weights considering values from $1$ to $5$ and found that the selected combination of weights yield the best PSNR.
@@
 Figure S17: Examples of HDR images reconstructed with our method and recent state-of-the-art.
-Figure S18: Examples of HDR images reconstructed with our method and recent state-of-the-art.